OpenStudy (anonymous):

Which is an equivalent form of the following? Help with this question.. square root of minus 32 y, end root, plus, square root of minus 98 y, end root

OpenStudy (anonymous):

\[\sqrt{-32y}+\sqrt{-98y}\]

OpenStudy (anonymous):

sorry thats the right question, so I got this as my answer. \[i \sqrt{130}\]

OpenStudy (anonymous):

\[\sqrt{-32y}+\sqrt{-98y}=\sqrt{\iota^232y}+\sqrt{\iota^298y}=\iota \sqrt{16\times2y}+\iota \sqrt{49\times2y}\] \[=4\sqrt{2y}\iota+7\sqrt{2y}\iota=11\sqrt{2y}~\iota\]

OpenStudy (anonymous):

oh why is it i^2 instead of just i

OpenStudy (anonymous):

-32=32*-1=32*iota^2

OpenStudy (anonymous):

\[\iota^2=-1\]